A river 200m wide is flowing at A rate of 4m/s.A boat crosses the river in minimum time . if velocity of the boat in still water is 8m/s then how far from the point directly opposite to the starting point boat reach

Grade:11

1 Answers

Kapil Khare
5 years ago
The boat will reach the other side in shortest time when it will have its velocity towards the other side that is it will have its velocity perpendicular to the the direction of flow of river. But as the river is also flowing the boat will drift in the direction of flow of river.
Width of the river = 200m
Velocity of boat in still water = 8m/s
So, time taken by boat to reach the other side of the river = (200)/(8) = 25s
Velocity of water = 4m/s
So, drift = 4*25 = 100m
So, boat will reach 100m far from the point directly opposite to the starting point.